MATH 375 – Probability and Statistics 1 A Short Questionnaire

Explanation

For use in some examples of data analysis later in the semester, please answer the questions below. “Home” refers to your permanent home address (not your HC residence during the school year, assuming you live on campus). Try to answer as accurately as you can. However, if some question is ambiguous or difficult to answer, you can leave it blank. Do not include your name on this sheet. Questions

1. What is your height in inches? (For instance, 5 feet 6 inches = 66 inches.)

2. What is your gender? (M or F)

3. What is your birthday? (month/day)

4. What is the zip code of your home address?

5. How many brothers and sisters (total) do you have?

6. How many first cousins do you have? Note: a first cousin is a child of a sibling (brother or sister) of one of your parents. Strictly speaking, only biological children of siblings of one of your biological parents “count.”

7. How many television sets are there in your home?

8. How many computers are there in your home? (Include both desktops and laptops, even if the laptops don’t spend all of their time there.)

9. How many pets are there in your home?

10. How many hours did you spend on Facebook in the past week? (nearest whole number of hours)